OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3)

OpenStudy (saifoo.khan):

Slope equals: 3-5 --- -2-1

OpenStudy (saifoo.khan):

2 --- 3

OpenStudy (saifoo.khan):

finding y intercpt: 3 = 2(-2)/3 + c 3 - 4/3 + c c = 5/3

OpenStudy (vijay):

3y = 2x +13

OpenStudy (saifoo.khan):

y = mx + b 3y = 2x + 13

OpenStudy (saifoo.khan):

Vijay is correct, i didnt took the LCm lol.